number 65&69 please
f(x) > 0 on (-0,1); f(x) < 0 on (1, 0) I c In Problems 53–74, summarize the pertinent information obtained by applying the
Answer #1

4 2 (0.0) (1.5.0) -2 0 (1.25, -1.526) -24697 f(x) = 0.52 -0.58 +4 e It has no roots function is always tive ou au at se=0 y=5 Range is [4,00) 1320 1320,0) f(x) = poi-10 (0,5) (1.386, 4) O 5